OpenStudy (anonymous):

(2-i)+(3i+2) help solve please

terenzreignz (terenzreignz):

Complex numbers? The constant i behaves like most any other variable. The only difference is that \(i^2\) is something special, in that it's equal to -1.

OpenStudy (anonymous):

(2-i)+(3i+2) =2-i + 3i+2= 4-2i